A New receiver just came out on the market, claiming it can deliver pass to pass accuracy of 5 inches at CEP.

What is the accuracy based on RMS in inches?
Answer #1

Circular Error Probability (CEP)  
CEP refers to the radius of a circle in which 50% of the values occur, i.e. if a CEP of 5
meters is quoted then 50% of horizontal point positions should be within 5 meters of the
true position. The radius of the 95% is often quoted and the term R95 used. R95 is CEP
with the radius of the 95% probability circle.

Distance Root Mean Squared (DRMS)  

DRMS is a single number that expresses 2D accuracy. The square root of the average of
the squared horizontal position errors.

Different ACccuracy of GPS

Accuracy Measures Formul:a Probability Definition The square root of the average of the squared horizontal positioin DRMS eirors. I 2νο; +o; | | Twice then Ris of the 2DRMS 95% 0.62o, +0.560% (Accurate when σ, / σ. > 0.3) R(0.62ay +0.567,) (R-2.08, when horizontal position errors The radius of circle centered at the true position, containing the position estimate with robability of 50% The radius of circle centered at the true position, containing the position estimate with probability of 95% CEP 5000 R95 95%

Accuracy Multipliers

CEP = 5 inches

The different accuracy based on RMS in inches is converted using above table.

2DRMS= 5*2.4= 12 inches

RMS(3D)=5*2.5= 12.5 inches

RMS(Horizontal)= 5*1.2= 6 inches